# Birgir Hrafnkelsson Morris Riedel Egill Skúlason

### v1

@@ 1.3 Overview
The introductory Chapter 2 describes basic preliminaries and pointers to the wider literature for the three fields under consideration, assumed to understood in subsequent chapters. Across the programming languages and academic domains covered, wall time efficiency forms the core constraint, starting from an efficient formulation of a finite element method for relativistic calculations in Fortran introduced in Chapter 3. Software development, and methods developed for visualization efficient double ended saddle searches are introduced together with the EON software suite, split across C++ and Python, with a raft of additional supporting libraries in Chapter 4.3, which also contains details on workflows and the further evolution of the software stack. Notably, we demonstrate how the scalable designs enable hiterto hard-to-implement workflows including the state of the art nudged elastic band methods and a novel hybrid method, the minimum mode following nudged elastic band (MMF-NEB). These then metastize into an implementation of a Gaussian Process acceleration for single ended saddle searches in C++ which explores a large, and albeit flawed benchmark set of calculations. The methodology is extended to surface calculations and vetted for use within proximal reaction network explorations. Exhaustive case-studies do not scale, so we explore performance modeling in high throughput regimes using Bayesian hierarchical methods in R and Stan to highlight computational bottlenecks in Chapter 6, before directly tackling data efficiency of Gaussian Process methods in Chapter 7. These efforts culminate in Chapter 8 with the optimal transport Gaussian process dimer (OTGPD), a framework leveraging optimal transport distance metrics which are chemically intuitive and transferable across systems. The OTGPD is shown to have state of the art wall time efficiency. Brief summary and conclusions pave the way for the reader to engage with the primary articles in the appendix 222on ArXiv these are not appended, but are freely available.
@@ 2.1.1 Rotational Step
The primary goal of the rotational step is to align the dimer’s orientation vector, $\hat{\mathbf{N}}$, with the minimum mode at the midpoint $\mathbf{R}$. This is achieved by rotating the dimer to minimize its total energy, $E=E_{1}+E_{2}$. The forces at the endpoints, $\mathbf{F}_{1}$ and $\mathbf{F}_{2}$, provide the necessary information to estimate the local curvature without computing the Hessian. The curvature $C$ along the dimer axis can be approximated using a finite difference:
@@
An effective rotational force, or torque, is derived from the atomic forces and used to drive the rotation, as depicted in Figure 6. This process is repeated using an optimization algorithm, such as conjugate gradient (CG), until the torque vanishes and $\hat{\mathbf{N}}$ is aligned with the minimum mode.
@@ 2.2 The Nudged Elastic Band (NEB) Method
Minimum mode following methods are useful for reaction network generation with applications to kinetics, however, when the end-points of a reaction are known, i.e. for a given reactant and product, often would like to determine likely paths between them in configuration space. A discrete approximation to a trajectory between the reactant and product can be formed through a “chain of states”. Conceptually, this can be viewed as an extension of the dimer images, and the subsequent string is then connected through springs.
@@ 2.3 Gaussian Process Regression
We begin by positing that the unknown Potential Energy Surface, a function $f(\mathbf{x})$, represents a single realization from a Gaussian Process. A Gaussian Process defines a probability distribution over a space of functions. Any finite collection of function values drawn from this process follows a joint multivariate normal (MVN) distribution.
@@
Each draw from the Gaussian Process forms a concrete MVN fully specified by a mean vector $\mathbf{m}$ and a covariance matrix $\mathbf{K}$. We construct these from a mean function $m(\mathbf{x})$, assumed to be zero throughout this work, and a covariance function or kernel $k(\mathbf{x},\mathbf{x}^{\prime})$ which determines the covariance between the energy and forces at different atomic configurations $\mathbf{x}$ and $\mathbf{x}^{\prime}$. The choice of kernel encodes prior assumptions about the functional form of the Potential Energy Surface. For systems with strong repulsive forces at short distances, the most commonly used kernel, the Squared Exponential (SE) kernel can be too restrictive [35].
We therefore employ an SE kernel based on an inverse distance metric [35, 36]. This provides a strong physical prior, leveling out the sharp increase in repulsive force when pairs of atoms get too close. This choice improves the model’s ability to learn high-gradient regions of the potential energy surface.
@@
In other words, we model the joint distribution of the energy and forces at multiple configurations $\{\mathbf{r}_{1},\mathbf{r}_{2},\dots,\mathbf{r}_{M}\}$ as a multivariate Gaussian distribution:
where:
  • $\mathbf{o}_{i}=\mathbf{o}(\mathbf{r}_{i})$ is the energy and force vector at configuration $\mathbf{r}_{i}$.
  • $\mathbf{K}$ is the covariance matrix. The elements of $\mathbf{K}$ are given by a kernel function $k(\mathbf{r}_{i},\mathbf{r}_{j})$ that measures the similarity between configurations $\mathbf{r}_{i}$ and $\mathbf{r}_{j}$. This kernel operates on the input space of atomic geometries ($\mathbb{R}^{N\times 3}$) and outputs the covariance between the combined energy and force vectors ($\mathbb{R}^{3N+1}$) at those geometries.
@@
where $\sigma_{f}^{2}$ is the signal variance, $\sigma_{c}^{2}$ is a constant offset, and $l_{\phi(i,j)}$ is the characteristic length scale for a specific pair type of atoms $\phi(i,j)$ 444only pairs are considered, a single atom type, O, for instance will not have an O-O term.
@@
The kernel hyperparameters $\boldsymbol{\theta}=\{\sigma_{f}^{2},\sigma_{c}^{2},l_{\psi},\dots\}$ are not known a-priori. They are learned from the training data by maximizing the log marginal log-likelihood (MLL):
@@ 2.4 Gaussian Process as an accelerator
The GPDimer method constructs a local surrogate model for the Potential Energy Surface through a finite, targeted set of samples. The approach functions like a reinforcement learning agent [43]. At each step of a search, the model decides whether to trust its current surrogate Potential Energy Surface or to query the “true” Potential Energy Surface via an expensive quantum chemical calculation. This “on-the-fly” model building refines a highly local and task-specific Potential Energy Surface with a minimal number of data points 555we will demonstrate results typically around 30 samples [36, 45].
@@ 3 Electronic structure calculations
To explore the Potential Energy Surface with the methods above, we must compute energies and atomic forces for given nuclear configurations. That task reduces to solving the many-electron Schrödinger equation in a form that scales to useful systems. We proceed by choice of representation: begin with a mean field to obtain a tractable one-body problem; then refine, when necessary, the physics or numerics.
@@ 3.1 Mean-field quantum chemistry
The simplest mean field theory approximation is the Hartree approximation, and involves assuming each electron moves in the Coulomb field generated by nuclei and the spherically averaged density of all electrons [52]. This local, multiplicative field removes two-body integrals and invites efficient solvers. It also leaves self-interaction uncorrected. An improvement on this, the Hartee-Fock method [53] treats electron correlation through an antisymmetrized wave function characterized with a single Slater determinant to incorporate exact exchange. As such, the non-local Fock operator introduces an integral operator which couples space points and removes self-interaction, though only exchange correlations are considered.
@@
with $\hat{h}=-\tfrac{1}{2}\nabla^{2}+v(\mathbf{r})$. In a spherical atom, orbitals separate, $P_{nl}(r)=rR_{nl}(r)$, and the Hartree term reduces to a 1D Poisson solve for $V_{H}(r)$. Exchange remains nonlocal after angular reduction. A multipole expansion yields
@@
From this basic formulation, subsequent expansions involving the occupancy of spin-orbitals yields unrestricted, restricted, and restricted open-shell Hartree-Fock methods [54], all of which are solved iteratively, through the so called “self-consistent field” approach. The addition of spin orbit occupancies at this stage does not change the fundamental nature of the mean field approximation. Since instantentious interactions are typically more repulsive than an average interaction, the difference between the exact energy and the self consistent field (SCF) approximation is the “correlation” energy, which can be approximated through more sophisticated treatments of the two-electron interaction, and lead to the “post-Hartree-Fock” methods [55]. These prescriptions are compelling mathematically, but in practice scale too poorly to be computed often for most systems. An numerically efficient alternative comes from a different angle, grounded in the density of electrons rather than exhaustively enumerating their positions.

### v2

@@ 1.2 Motivation
1.3 Research objectives and hypotheses
The central thesis of this work is that the efficiency of rare-event sampling is constrained not just by the cost of electronic structure, but by the rigidity of standard software architectures and the fragility of optimization algorithms. To address this towards the in-silico control of materials to accelerate scientific applications, we investigate the following across various length and time scales:
Relativistic spurious state evasion
Can a squared Hamiltonian formulation in a finite element modular Fortran framework provide state of the art stability222by avoiding spurious states since no variational principle holds for the Dirac case and numerical performance for single atom all-electron mean-field calculations of both the Schrodinger and Dirac equations?
Saddle diagnostics
Can the interpretation of saddle structures be improved beyond the “eyeball norm” and baseline methods of calculating bonds based on covalent radii with approximate energy surfaces? Can optimization trajectories towards these points be improved?
Software designs
Does the modernization of existing software unlock novel scientific algorithms for reducing the number of samples needed to find a saddle point?
Internal coordinates
Can performance gains from complex internal coordinate generating / coordinate drive methods be beaten by better engineered Cartesian coordinate based local surrogate saddle searches?
Statistical performance modeling
Average-cost benchmarks and point estimates of performance mask algorithmic reliability despite havnig poor statistical backing due to repeated data. Can more robust Bayesian methods lead to novel insights?
Wall-time efficiency
GP acceleration has been known to be theoretically efficient for almost a decade, but no implementation has managed to show high reliability and wall-time gains in time, can this be alleviated by careful analysis of hyperparameters and software?
@@ 1.4 Overview
The introductory Chapter 2 outlines basic preliminaries and provides pointers to the wider literature for the three fields under consideration. Across the programming languages and academic domains covered, wall-time efficiency functions as the core constraint, as illustrated in Figure 5.
We begin with the fundamental physics in Chapter 3, which introduces featom, a high-order finite element method (FEM) solver for the Dirac and Schroedinger equations in Fortran. Unlike the frozen core approximations common to valence chemistry, featom facilitates heavy element relativistic calculations in the all-electron mean-field approximation. This tool stands distinct from the kinetic methods discussed in the remainder of the thesis.
Figure 5: The structural connectivity of this dissertation. Wall-time efficiency serves as the primary constraint driving innovation across three domains: relativistic electronic structure (cyan), chemical kinetics software (purple), and statistical diagnostics (amber). EON (Chapter 4) functions as the central orchestrator, managing interfaces to potential energy and force calculators and maintaining a tight, bidirectional coupling with the distinct GP-accelerated package developed for the GPDimer in Chapter 5 and the OTGPD in 7. Quantification via the Bayesian performance metrics in Chapter 6 and visual diagnostics in Chapter 4 round out the software and methods developed.
Chapter 4.3 shifts focus to the EON software suite, a hybrid C++ and Python framework designed for long-timescale simulations. We detail the client-server architecture and the evolution of the software stack. Notably, we demonstrate how these scalable designs enable hitherto hard-to-implement workflows, including state-of-the-art nudged elastic band methods and a novel hybrid method, the minimum mode following nudged elastic band (MMF-NEB).
These efforts evolve into an implementation of Gaussian Process acceleration for single-ended saddle searches in C++, exploring a large, though historically challenging, benchmark set. We extend the methodology to surface calculations and vet it for use within proximal reaction network explorations. Because exhaustive case studies fail to scale indefinitely, Chapter 6 explores performance modeling in high-throughput regimes. We employ Bayesian hierarchical methods in R and Stan to highlight computational bottlenecks, before directly tackling the data efficiency of Gaussian Process methods in Chapter 7.
These efforts culminate in Chapter 8 with the OTGPD, a framework leveraging optimal transport distance metrics. This approach proves chemically intuitive and transferable across systems, demonstrating state-of-the-art wall-time efficiency. A brief summary and conclusions pave the way for the reader to engage with the primary articles in the appendix 333on ArXiv these are not appended, but are freely available.
@@ 2.1.1 Rotational Step
The primary goal of the rotational step involves aligning the dimer orientation vector, $\hat{\mathbf{N}}$, with the minimum mode at the midpoint $\mathbf{R}$. To this end, we rotate the dimer to minimize the total energy, $E=E_{1}+E_{2}$. With this, an effective rotational force, or torque $\mathbf{F}_{rot}$ can be derived from the forces at the endpoints, $\mathbf{F}_{1}$ and $\mathbf{F}_{2}$, to estimate the change in local curvature without computing the Hessian. The curvature $C$ along the dimer axis can be approximated using a finite difference:
@@
This $\mathbf{F}_{rot}$ defines the search direction perpendicular to $\hat{\mathbf{N}}$, depicted in Figure 7. We use an optimization algorithm, such as conjugate gradient (CG), to determine the orientation at which $\hat{\mathbf{N}}$ is aligned with the minimum mode. Practically, we recognize that the curvature within a rotational plane spanned by the current orientation $\hat{\mathbf{N}}$ and the normalized torque direction approximates a sinusoidal curve
$C(\phi)=A+B\cos(2\phi)+D\sin(2\phi)$ (4)
By evaluating the gradients at the current position we determine the angle which minimizes the curvature along a given search direction, and the dimer rotates to this orientation. This process occurs without the coordinates of the midpoint of the dimer changing and continues until the predicted rotation angle falls below a fixed convergence threshold. For standard calculations a rather loose tolerance of 5 degrees typically suffices.
@@ 2.2 The Nudged Elastic Band (NEB) Method
Minimum mode following methods prove useful for reaction network generation and general kinetic applications. However, when we distinguish specific reactant and product basins, the objective shifts. In such cases, we seek to determine the most probable paths connecting these fixed points in configuration space. A discrete approximation to a trajectory between the reactant and product can be formed through a “chain of states”. Conceptually, this can be viewed as an extension of the dimer images, and the subsequent string is then connected through springs.
@@ 2.2.6 Optimization and Path Analysis
2.3 Acceleration strategies
The core methodology described thus-far has formed an integral part of computational chemistry calculations for several decades. In order to obtain results on larger systems or longer time-scales, these must be accelerated. Within this thesis we consider two primary approaches to this acceleration:
Software design
As an engineering endeavor, generating efficient machine code leads to faster solutions and thus unlocks larger systems. Under this broad remit falls parallelization strategies, caching, handling large data, better use of specialized hardware like graphics processing unit (GPU) computations etc.
Algorithmic improvements
The total computational time depends on the algorithms employed, which in turn demand energy and atomic force evaluations from high-level theory. Consequently, algorithms that minimize the number of calculations required to achieve a specific solution accuracy constitute the primary metric for optimization. Utilizing cheaper approximate surfaces for portions of the algorithm represents a significant focus of current research.
An ideal surrogate demands ease of computation, autonomy from manual intervention, and “transparent” progression without the manual construction of training sets. We therefore mandate that the method:
Learns on the fly
Efficiently learning from low amounts of often correlated data.
Encode physics
Providing, at minimum, a functional form that aids analysis.
Trains quickly
Enabling the efficient incorporation of new points into the model.
We explicitly define the scope: strictly avoiding the goal of matching the underlying potential energy surface globally. These surrogates serve only to guide algorithms toward specific points in the solution space; thus, absolute errors in energy and forces hold little value compared to local gradient accuracy. Similarly, we exclude general energy surface construction, the basis of machine learning interatomic potentials. We therefore focus on regression techniques [11]. Computing atomic descriptors are based on cartesian coordinates [34] requires both user intervention for parameter selection, and increases the computational cost of on-the-fly learning. Consequently, an ideal surrogate for our consideration operates directly on raw Cartesian coordinates. We briefly evaluate the candidates:
Linear regression
The oldest fitting technique, which may be competitive with enough data over a small region of space, but only with descriptors to account for non-linearities, which are often incomplete by construction [35].
Neural networks
Deep neural networks [36] interpolate arbitrary systems well and can encode specific qualities through filters (e.g., convolutional architectures). However, they tend to overfit in the low-data regime [37], and require careful selection of both data and long training times for the hyperparameters. The resulting models have no closed form analytical solution and rely on being able to backpropagate gradients [38, 39], which also forms a computational bottleneck.
Gaussian processes
Often understood as limiting cases of spline models [40], neural networks [41, 42], quadrature rules [43], stochastic partial differential equations [44], finite realizations of these methods form multivariate normal distributions [45] which are analytically and computationally compact. The connection to standard statistics enables a rich set of uncertainity quantification, and the ability to constrain the function spaces based on distance measures between data points makes for more interpretable surrogates.
With this in mind, we move on with further contextualizing the Gaussian Process for molecular systems, though we note that descriptor based methods have also been applied for activated searches [46].
@@ 2.4 Gaussian Process Regression
We begin by positing that the unknown Potential Energy Surface, a function $f(\mathbf{x})$, represents a single realization from a Gaussian Process. A Gaussian Process defines a probability distribution over a space of functions. Any finite collection of function values drawn from this process follows a joint multivariate normal (MVN) distribution. It is worth noting that such an assumption can be justified rather rigorously for molecular potential energy surfaces with pronounced global minimum as demonstrated by [47] in terms of an asymptotic study of vibrational degrees of freedom, which leads to the potential energy surface on a random coordinate frame manifesting as a sum of many contributions, which in turn through the Central limit theorem [11] leads to an MVN perspective, equivalent to a perturbative approach for small molecules.
Such theoretical modeling of molecular potential energy surfaces [47] rests upon three postulates: the potential constants represent averages over all unitary transformations; the topology possesses a distinct global minimum; and the number of vibrational modes approaches the asymptotic limit. While these constraints, in particular the topological requirement strictly preclude the multiple-minima landscapes inherent to the reactive systems under investigation here, the derivation nevertheless offers robust physical support for approximating the energy surface locally as a multivariate Gaussian.
@@
Each draw from the Gaussian Process forms a concrete MVN fully specified by a mean vector $\mathbf{m}$ and a covariance matrix $\mathbf{K}$. We construct these from a mean function $m(\mathbf{x})$, assumed to be zero throughout this work, and a covariance function or kernel $k(\mathbf{x},\mathbf{x}^{\prime})$ which determines the covariance between the energy and forces at different atomic configurations $\mathbf{x}$ and $\mathbf{x}^{\prime}$. The choice of kernel encodes prior assumptions about the functional form of the Potential Energy Surface. For systems with strong repulsive forces at short distances, the most commonly used kernel, the infinitely differentiable squared exponential (SE) kernel can be too restrictive [49], which is only partially resolved by choosing different kernels like the Matern. The stationarity of the SE kernel, i.e. the assumption of uniform fluctuations across the domain makes it difficult to handle the steep gradients of repulsive walls in Cartesian space. By transforming the input space through an inverse distance metric, we effectively “precondition” the energy surface by homogenizing the effective length scale of the landscape, which enables the Gaussian Process to resolve high gradient regions without sacrificing resolution in energetic wells. We therefore employ a SE kernel based on an inverse distance metric [49, 50] since it provides a strong physical prior, leveling out the sharp increase in repulsive force when pairs of atoms get too close.
@@
In other words, we model the joint distribution of the energy and forces at multiple configurations $\{\mathbf{r}_{1},\mathbf{r}_{2},\dots,\mathbf{r}_{M}\}$ as a multivariate Gaussian distribution Eq. LABEL:eq:mvn_pred and we now identify the elements of $\mathbf{K}$ are given by a kernel function $k(\mathbf{r}_{i},\mathbf{r}_{j})$ that measures the similarity between configurations $\mathbf{r}_{i}$ and $\mathbf{r}_{j}$. This kernel operates on the input space of atomic geometries ($\mathbb{R}^{N\times 3}$) and outputs the covariance between the combined energy and force vectors ($\mathbb{R}^{3N+1}$) at those geometries.
@@
where $\sigma_{f}^{2}$ is the signal variance, $\sigma_{c}^{2}$ is a constant offset, and $l_{\phi(i,j)}$ is the characteristic length scale for a specific pair type of atoms $\phi(i,j)$ 555only pairs are considered, a single atom type, O, for instance will not have an O-O term.
@@
The kernel hyperparameters $\boldsymbol{\theta}=\{\sigma_{f}^{2},\sigma_{c}^{2},l_{\psi},\dots\}$ are not known a-priori. Most commonly, the hyperparameters derive from the marginal likelihood. Practically, optimizing takes place in log-space ($\eta=\log\theta$) since this provides better numerical scaling through two benefits.
Firstly, it implicitly maintains the strict positivity constraints required by variances and length scales. The mapping $\theta=\exp(\eta)$ ensures that the optimization algorithm, which typically operates in unconstrained Euclidean space, yields strictly positive physical parameters without requiring barrier functions or Lagrange multipliers.
Secondly, logarithmic scaling preconditions the objective function landscape. In chemical physics applications, hyperparameters frequently vary by orders of magnitude. A gradient step size appropriate for a large signal variance $\sigma_{f}^{2}$ could prove disastrously large for a short-range length scale $l_{\psi}$. Working in the log-space equilibrates these sensitivities, as a constant step size in $\boldsymbol{\eta}$ corresponds to a proportional change in $\boldsymbol{\theta}$ rather than an absolute one. This effectively balances the gradient contributions across disparate dimensions and accelerates convergence.
From a practical standpoint, the gradient computation for the transformed parameters follows the chain rule:
$\frac{\partial\mathcal{L}}{\partial\eta_{i}}=\frac{\partial\mathcal{L}}{\partial\theta_{i}}\frac{\partial\theta_{i}}{\partial\eta_{i}}=\frac{\partial\mathcal{L}}{\partial\theta_{i}}\theta_{i}$ (29)
This scaling factor $\theta_{i}$ naturally dampens the gradient for small parameters and amplifies it for large ones, further stabilizing the descent. Thus the kernel hyperparameters are learned from the training data by maximizing the logarithm of the marginal log-likelihood (MLL):
@@ 2.4.1 Scaling in Time and Storage
Inference time
Once the kernel components are determined, prediction involves a single kernel vector between the new point and training points, so this scales linearly with the number of training points.
@@ 2.5 Gaussian Process as an accelerator
The GPDimer method constructs a local surrogate model for the Potential Energy Surface through a finite, targeted set of samples. The approach functions like a reinforcement learning agent [56]. At each step of a search, the model decides whether to trust its current surrogate Potential Energy Surface or to query the “true” Potential Energy Surface via an expensive quantum chemical calculation. This “on-the-fly” model building refines a highly local and task-specific Potential Energy Surface with a minimal number of data points 666we will demonstrate results typically around 30 samples [50, 58].
@@
2.6 Conclusions
This chapter has delineated the theoretical and algorithmic framework necessary for navigating high-dimensional potential energy surfaces. We detailed the topological search methods based on the number of known minima, specifically the single ended Dimer and double ended Nudged Elastic Band algorithms which serve as the primary vehicles for locating first-order saddle points and minimum energy pathways. While these techniques provide robust convergence properties, in practice, their reliance on expensive ab-initio force evaluations necessitates an acceleration strategy capable of operating within the strict data constraints of on-the-fly exploration.
Our analysis identified Gaussian Process Regression as the optimal surrogate for this regime. By combining the asymptotic derivation of local energy landscapes with a physically motivated inverse-distance kernel, we show how one constructs a model that balances flexibility with the regularization required by repulsive atomic cores. Crucially, this approach bypasses the need for the extensive datasets required by global machine learning potentials, favoring instead a rigorous local approximation that respects the underlying physics of the reaction channel. With these mathematical and computational foundations established, we later proceed to their practical implementation and application in the study of complex reactive systems.
However, a surrogate model acts merely as an efficient interpolator; it possesses no intrinsic knowledge of the quantum mechanics. The fidelity of any reaction path search depends ultimately on the accuracy of the underlying electronic structure calculations. Furthermore, while we discuss methods to minimize the number of evaluations, the computational cost per evaluation remains a critical bottleneck, particularly for heavy-element systems where relativistic effects dominate. Therefore, before applying these exploration algorithms to complex reactive systems, we cover some eletronic structure theory for chemical systems and foray into a rigorous and efficient foundation for computing the potential energy surface for single atom systems. The subsequent chapter addresses this challenge, introducing a high-order finite element approach to the self-consistent field problem that balances relativistic accuracy with numerical efficiency.
@@ 3 Electronic structure calculations
To explore the Potential Energy Surface with the methods described previously, we must compute energies and atomic forces for specific nuclear configurations. This task reduces to solving the many-electron Schrödinger equation in a form that scales to useful systems. We proceed by choice of representation: we begin with a mean field to obtain a tractable one-body problem; then we refine the physics or numerics as necessary.
After a brief introduction to the general electronic structure methods for chemical systems, we focus on a state of the art solver, featom which solves for isolated single atom systems, the Dirac and Schrödinger equations.
@@ 3.1 Mean-field quantum chemistry
The Hartree approximation represents the simplest mean field theory approximation. It assumes that each electron moves in the Coulomb field generated by nuclei and the spherically averaged density of all other electrons [65]. This local, multiplicative field removes two-body integrals and invites efficient solvers. However, it leaves self-interaction uncorrected. The Hartree-Fock (HF) method [66] improves upon this by treating electron correlation through an antisymmetrized wave function characterized by a single Slater determinant to incorporate exact exchange. As such, the non-local Fock operator introduces an integral operator which couples space points and removes self-interaction, though it considers only exchange correlations.
@@
with $\hat{h}=-\tfrac{1}{2}\nabla^{2}+v(\mathbf{r})$. In a spherical atom, orbitals separate as $P_{nl}(r)=rR_{nl}(r)$, and the Hartree term reduces to a 1D Poisson solve for $V_{H}(r)$. Exchange remains nonlocal after angular reduction. A multipole expansion yields
@@
From this basic formulation, subsequent expansions involving the occupancy of spin-orbitals yield unrestricted, restricted, and restricted open-shell Hartree-Fock methods [67]. All of these are solved iteratively through the self consistent field (SCF) approach. The addition of spin-orbit occupancies at this stage leaves the fundamental nature of the mean field approximation unchanged. Since instantaneous interactions typically prove more repulsive than an average interaction, the difference between the exact energy and the SCF approximation constitutes the “correlation” energy. More sophisticated treatments of the two-electron interaction can approximate this energy, leading to “post-Hartree-Fock” methods [68]. These prescriptions offer mathematical rigor but scale too poorly for frequent computation in large systems. A numerically efficient alternative arrives from a different angle: grounding the theory in the density of electrons rather than exhaustively enumerating their positions.